# Estimating Player Performance in Different Contexts Using Fine-tuned Large Events Models

### v1

@@
Abstract.
This paper introduces an innovative application of Large Event Models (LEMs), akin to Large Language Models, to the domain of soccer analytics. By learning the ”language” of soccer — predicting variables for subsequent events rather than words — LEMs facilitate the simulation of matches and offer various applications, including player performance prediction across different team contexts. We focus on fine-tuning LEMs with the WyScout dataset for the 2017-2018 Premier League season to derive specific insights into player contributions and team strategies. Our methodology involves adapting these models to reflect the nuanced dynamics of soccer, enabling the evaluation of hypothetical transfers. Our findings confirm the effectiveness and limitations of LEMs in soccer analytics, highlighting the model’s capability to forecast teams’ expected standings and explore high-profile scenarios, such as the potential effects of transferring Cristiano Ronaldo or Lionel Messi to different teams in the Premier League. This analysis underscores the importance of context in evaluating player quality. While general metrics may suggest significant differences between players, contextual analyses reveal narrower gaps in performance within specific team frameworks.
Large Event Models, Sports Analytics, Predictive Modeling, Fine-tuning, Deep Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networksccs: Information systems Data miningccs: Computing methodologies Discrete-event simulation
Figure 1. A comparison between the expected points distribution of the record-breaking Manchester City 2017/18 team without and with the world’s two biggest soccer superstars, Cristiano Ronaldo and Lionel Messi
1. Introduction
Effectively navigating decision-making in the sports domain, especially when substantial financial investments in players or managers are at stake, presents a complex challenge. Despite the substantial growth in using data to improve decision-making in recent decades (Statsbomb, [n. d.]), its application often needs a comprehensive approach. An illustrative example highlighting this need is the assessment of compatibility between a specific player and a particular team. While various companies in the market offer ”AI-based solutions” for determining this compatibility, there is a notable absence of published validation studies for such solutions. Many of these solutions may be grounded in a linear paradigm, whereas the complexity of the problem clearly suggests the requirement for a non-linear method.
In the meantime, roughly 50% of soccer transfers fail (Tom Worville, 2021), which can be linked to diverse factors (see Table 1). Certain aspects, such as adaptability issues, might be beyond the control of sporting directors or other decision-makers. We assert that it is possible to quantify whether a specific player aligns with a particular playing style and if his integration into a team ultimately proves beneficial, but this cannot be achieved through traditional methods.
@@ 1. Introduction
Table 1. Why transfers fail according to Ian Graham, compiled by Tom Worville (Tom Worville, 2021). *Reasons that can be addressed in our framework.
In a broad sense, this study aims to evaluate a player’s impact on a soccer team, thus aligning with various existing studies. Numerous methods employ frameworks that assess events based on game state value, commonly called possession value (Statsbomb, [n. d.]). One example is the Valuing Actions by Estimating Probabilities (VAEP) model, which treats the estimation of game state value as a machine learning challenge (Decroos et al., 2019). It calculates the probability of an action leading to a goal within subsequent actions. Another approach involves the expected threat (xT) metric, assigning values to different areas of the pitch based on the likelihood of scoring from those zones (Statsbomb, [n. d.]). These methods go beyond event data, providing solutions for tracking data (Fernández et al., 2021). Furthermore, we can also see some relationship between our approach and models designed to estimate player transfer fees (McHale and Holmes, 2023), valuing players over extended periods (Mendes-Neves et al., 2022), and other related aspects. The significance of these methods cannot be overstated. Research indicates that following the Moneyball paradigm (Lewis, 2004), most teams rapidly adopted this data-centric approach in their recruitment strategies (Hakes and Sauer, 2006).
However, the current approaches neglect to account for the impact of contextual factors on the values they measure. This gap becomes evident when addressing a fundamental question: How would these values change if a player were to join a different team? Despite notable research on how VAEP varies with a player’s league (The Come On Man, [n. d.]) and the influence of players on action outcomes (Tahmeed Tureen and SBH Olthof, 2022), this particular question remains largely unanswered.
@@
This methodology addresses a specific issue: while LLMs excel in creative tasks, they struggle in reasoning tasks (Huang and Chang, 2023). For example, LLMs face challenges in generating a concise list of players who could improve a team. LEMs, on the other hand, enable the simulation of diverse contexts and the potential behavior of a player within those contexts. This capability facilitates informed decision-making based on multiple Key Performance Indicators (KPIs). These include anticipated contributions to the team in terms of points, shots, key passes, crosses, and set-piece goals.
To answer these questions, we fine-tune LEMs to comprehend team and player behaviors across different scenarios. Our analysis involves assessing teams from the 2017/18 English Premier League (EPL) season, demonstrating LEMs’ capacity to capture distinct soccer team behaviors. Additionally, we conduct a comparative analysis for each Premier League team, evaluating the potential impact of acquiring either Messi or Ronaldo. This comparison helps clarify our methodology’s strengths and limitations, providing insights into its accuracy and areas of improvement.
@@
  • Section 2 provides the background knowledge on Large Events Models.
@@
2. Background
2.1. Large Events Models
@@
The Large Events Model (LEM) draws inspiration from Large Language Models (LLMs), applying their sequential prediction approach to the domain of soccer (Mendes-Neves et al., 2024). LLMs utilize the context provided by preceding words to forecast the subsequent word in a text, with each new word modifying the context for continuous, coherent text generation. Similarly, LEMs utilize the current state of a soccer game to anticipate the next event. This model aims to predict the forthcoming event in a series based on the current game state (i.e., current result, previous events). Leveraging the sequential nature of soccer event data, LEMs enable large-scale simulation of soccer matches. These simulations are grounded entirely in data, with the model perpetually updating the game state with each new predicted event, as depicted in Figure 2.
@@
Figure 2. LEMs and LLMs work on the same principle: given the initial context that contains all current information, it can forecast the next token. This token then updates the context iteratively until an exit criterion is met. For LLMs, the tokens are words. For LEMs, the tokens are events.
@@
This approach provides many advantages over traditional machine learning approaches. It can be used to infer basic statistical approaches like expected goals (Pollard et al., 2004) or events to simulate large amounts of data (Mendes-Neves et al., 2021). Most relevant is that it offers a framework that uses the same backbone to generate many insights through simulation. Even smaller LEMs can be used to simulate a restricted version of the soccer match, focusing only on key events (Simpson et al., 2022; Yeung et al., 2023).
In our case, the LEM we are going to use utilizes the following features from the previous event to provide the context: (1) the action type, (2) time, (3) location x and y, (4) Boolean variables indicating if the action is accurate or a goal, and (5) the current score of the game.
@@
Then, our LEM uses a phased approach to ensure that the multiple variables we need to forecast are aligned. For example, if the following action is a pass, then the likelihood of being an accurate action is much higher when compared to being a shot. If we tried to forecast the two variables simultaneously, the error would be much higher than forecasting them sequentially. Therefore, the LEM forecasts the variables in the following order:
  1. (1)
    The next action type.
  2. (2)
    The accuracy of the next action and if the next action is a goal.
  3. (3)
@@
    The next action location, time elapsed until the next action, and if the home team performs the next action.
Note that every forecast is probabilistic, which means that, for the same context, the output will vary according to the forecasted probabilities.
@@
LEMs have advantages and limitations. The positives include their foundational role in creating various metrics for soccer players, ranging from general performance in a season to specific impacts on aspects like shooting quality in particular zones. Simulating numerous games also allows for extracting patterns in the data, enabling comparison with baseline or affected patterns to create insightful metrics [15]. However, a significant limitation is that the model simulates generic football matches, as it was trained on data from all teams, resulting in average game simulations. In the next Section, we will propose a methodology to enable context-specific insights by fine-tuning the models to replicate the behavior of specific teams and players to understand how changing a player’s context affects their performance.
@@ 3.1. Data
This study relies on the Public Wyscout dataset (Pappalardo et al., 2019) as the primary source of football event data. This dataset encompasses fixtures from top-tier leagues, including the Premier League (England), La Liga (Spanish), Serie A (Italy), Bundesliga (Germany), and Ligue 1 (France), with a focus on the 2017/2018 season. It provides an extensive range of events, from passes and shots to aerial duels and saves. We opted for this dataset since it was the most extensive available at the start of the project. Although the size of the dataset limits the comprehensiveness of the study, it was important for us to provide an open platform to incentive collaboration for further developments.
@@
3.2. Fine Tuning LEMs to Learn Different Contexts
Fine-tuning means adjusting a pre-trained model on a specific task or dataset to improve performance. It involves further training a model already trained on a large dataset, typically on a more specialized or smaller dataset related to the task (Radford et al., 2018). The overall process can be seen in Figure 3.
Figure 3. This figure depicts the two-stage process of developing a fine-tuned LEM model: first, we use a large dataset to build a LEM, then we fine-tune it using specific data about our target.
To learn new contexts, we fine-tune the models to reflect the specific behaviors of the teams. In essence, we change the context-agnostic LEMs to perform another task. Instead of forecasting the next event in an average game of soccer, it will forecast the next event in a game within a predefined context. For this, we select the subset of data that is of interest. For example, if we want to build a model replicating the behavior of one team playing at home, we fine-tune the model using data from that specific team playing at home. Table 2 presents the different types of contexts that we can learn in LEMs.
LEM Fine-Tune Team Opponent Player Player
@@ 3.2. Fine Tuning LEMs to Learn Different Contexts
Table 2. The types of fine-tuning that we can perform on LEMs. The type ”Team” focuses on replicating a single team’s behavior against the league average. The type ”Player” focuses on estimating the impact of the player on the average team of the league. ”Player Addition” includes data from a new player in the context of the new team. ”Player Replacement” does the same but excludes data from the player being replaced. In the ”Team Face-off” type, we have the data of two teams: the home team and the away team.
This paper focuses on the first four types of fine-tuned LEMs. We have excluded the ”Team Face-off” model from this analysis. The rationale behind this exclusion is that ”Team Face-off” primarily centers on evaluating potential tactical outcomes of two opposing teams rather than providing insights into player recruitment, which is the primary focus of this study.
3.3. Player Addition vs Player Replacement
Two categories can be used to measure a player’s performance within a new context: ”Player Addition” and ”Player Replacement.” The difference between both approaches is that Player Replacement actively removes the data from a competing player in the squad. On top of adding the influence of the incoming player, we remove the existing influence of the previous player on our squad. Therefore, instead of the training process slowly overwriting the data from the players, it is removed, allowing the training process to start from a blank slate. Table 3 shows the players chosen to be replaced. The criterion for selection of these players was ”the attacking player with the least amount of game time from the top 11 most used players in the season”.
@@ 3.3. Player Addition vs Player Replacement
Table 3. The list of players that are replaced in the Player Replacement fine-tuning.
3.4. Parameter Tuning
@@
However, we recommend optimizing these parameters through Bayesian optimization methods (same as the original LEM). We did not incur on this process due to our computational constraints. We acknowledge that avoiding the parameter optimization step can induce errors and biases in our results. However, for this work and accounting for the remaining limitations of the methodology we propose, this error is negligible. Nonetheless, this step is essential to account for applications of this method.
(1) $\text{Batch Size}=\log(n^{2}),\quad\text{where}\quad 32\leq\text{Batch Size}\leq 256$
The definition of the batch size is the most critical parameter in this work. Fine-tuning a model requires finding the right balance between frequent updates through backpropagation and the training speed of the models (Radford et al., 2018). Iteratively, we found Equation 1 to provide the right balance for our use case. The other relevant parameter we must set is the number of simulations for each LEM. As visible in Figure 4, the larger the number of simulations, the lower the variance of the results. We opted for a value of 2500 simulations since it finds a reasonable compromise between simulation time and low variance in the results. Furthermore, we train 10 iterations for each model, allowing us to evaluate the variability of the training process.
Figure 4. This figure illustrates that the outcome variability originates from the model’s training and fine-tuning rather than the simulation process. At the threshold of 3000 simulations, a sequence of 10 consecutive wins is required to alter the expected points by a mere +0.01. The fluctuations seen in the results are attributable to the training and fine-tuning of the model, not the simulation process itself. Additionally, the figure shows that the error distribution follows a normal curve, indicating that, over an extended period, the average error in simulations is expected to converge towards zero.
3.5. Limitations

### v2

@@
Abstract
This paper introduces an innovative application of Large Event Models (LEMs), akin to Large Language Models, to the domain of soccer analytics. By learning the "language" of soccer — predicting variables for subsequent events rather than words — LEMs facilitate the simulation of matches and offer various applications, including player performance prediction across different team contexts. We focus on fine-tuning LEMs with the WyScout dataset for the 2017-2018 Premier League season to derive specific insights into player contributions and team strategies. Our methodology involves adapting these models to reflect the nuanced dynamics of soccer, enabling the evaluation of hypothetical transfers. Our findings confirm the effectiveness and limitations of LEMs in soccer analytics, highlighting the model’s capability to forecast teams’ expected standings and explore high-profile scenarios, such as the potential effects of transferring Cristiano Ronaldo or Lionel Messi to different teams in the Premier League. This analysis underscores the importance of context in evaluating player quality. While general metrics may suggest significant differences between players, contextual analyses reveal narrower gaps in performance within specific team frameworks.
[…]
2338
1 Introduction
Effectively navigating decision-making in the sports domain, especially when substantial financial investments in players or managers are at stake, presents a complex challenge. Despite the substantial growth in using data to improve decision-making in recent decades [14], its application often needs a comprehensive approach. An illustrative example highlighting this need is the assessment of compatibility between a specific player and a particular team. While various companies in the market offer "AI-based solutions" for determining this compatibility, there is a notable absence of published validation studies for such solutions. Many of these solutions may be grounded in a linear paradigm, whereas the complexity of the problem clearly suggests the requirement for a non-linear method.
In the meantime, roughly 50% of soccer transfers fail [17], which can be linked to diverse factors (see Table 1). Certain aspects, such as adaptability issues, might be beyond the control of sporting directors or other decision-makers. We assert that it is possible to quantify whether a specific player aligns with a particular playing style and if his integration into a team ultimately proves beneficial, but this cannot be achieved through traditional methods.
@@ 1 Introduction
Table 1: Why transfers fail according to Ian Graham, compiled by Tom Worville [17]. *Reasons that can be addressed in our framework.
In a broad sense, this study aims to evaluate a player’s impact on a soccer team, thus aligning with various existing studies. Numerous methods employ frameworks that assess events based on game state value, commonly called possession value [14]. One example is the Valuing Actions by Estimating Probabilities (VAEP) model, which treats the estimation of game state value as a machine learning challenge [1]. It calculates the probability of an action leading to a goal within subsequent actions. Another approach involves the expected threat (xT) metric, assigning values to different areas of the pitch based on the likelihood of scoring from those zones [14]. These methods go beyond event data, providing solutions for tracking data [2]. Furthermore, we can also see some relationship between our approach and models designed to estimate player transfer fees [6], valuing players over extended periods [8], and other related aspects. The significance of these methods cannot be overstated. Research indicates that following the Moneyball paradigm [5], most teams rapidly adopted this data-centric approach in their recruitment strategies [3].
However, the current approaches neglect to account for the impact of contextual factors on the values they measure. This gap becomes evident when addressing a fundamental question: How would these values change if a player were to join a different team? Despite notable research on how VAEP varies with a player’s league [16] and the influence of players on action outcomes [15], this particular question remains largely unanswered.
@@
This methodology addresses a specific issue: while LLMs excel in creative tasks, they struggle in reasoning tasks [4]. For example, LLMs face challenges in generating a concise list of players who could improve a team. LEMs, on the other hand, enable the simulation of diverse contexts and the potential behavior of a player within those contexts. This capability facilitates informed decision-making based on multiple Key Performance Indicators (KPIs). These include anticipated contributions to the team in terms of points, shots, key passes, crosses, and set-piece goals.
To answer these questions, we fine-tune LEMs to learn team and player behaviors across different scenarios. Our analysis involves assessing teams from the 2017/18 English Premier League (EPL) season, demonstrating LEMs’ capacity to capture distinct soccer team behaviors. Additionally, we conduct a comparative analysis for each Premier League team, evaluating the potential impact of acquiring either Messi or Ronaldo. This comparison helps clarify our methodology’s strengths and limitations, providing insights into its accuracy and areas of improvement.
@@
  • Section 2 provides the background knowledge on LEMs.
@@
2 Background
2.1 Large Events Models
@@
LEMs draw inspiration from LLMs, applying their sequential prediction approach to the domain of soccer [9]. LLMs utilize the context provided by preceding words to forecast the subsequent word in a text, with each new word modifying the context for continuous, coherent text generation. Similarly, LEMs utilize the current state of a soccer game to anticipate the next event. This model aims to predict the forthcoming event in a series based on the current game state (i.e., current result, previous events). Leveraging the sequential nature of soccer event data, LEMs enable large-scale simulation of soccer matches. These simulations are grounded entirely in data, with the model perpetually updating the game state with each new predicted event, as depicted in Figure 1.
@@
Figure 1: LEMs and LLMs work on the same principle: given the initial context that contains all current information, it can forecast the next token. This token then updates the context iteratively until an exit criterion is met. For LLMs, the tokens are words. For LEMs, the tokens are events.
@@
This approach provides many advantages over traditional machine learning approaches. It can be used to infer basic statistical approaches like expected goals [11] or to simulate large amounts of data [7]. Most relevant is that it offers a framework that uses the same backbone to generate many insights through simulation. Even smaller LEMs can be used to simulate a restricted version of the soccer match, focusing only on key events [13, 18].
@@
Our LEM uses a phased approach to ensure that the multiple variables we need to forecast are aligned. For example, if the following action is a pass, then the likelihood of being an accurate action is much higher when compared to being a shot. If we tried to forecast the two variables simultaneously, the error would be much higher than forecasting them sequentially. Therefore, the LEM sequentially forecasts the variables in the following order:
  1. 1.
    Model Type forecasts the next event type.
  2. 2.
    Model Accuracy forecasts the accuracy of the next event and if the next event is a goal.
  3. 3.
@@
    Model Data forecasts the next event location, the time elapsed until the next event, and if the home team performs the next event.
The LEM forecasts the probability distribution for each variable. For example, the output of Model Type is the probability for each possible event. In a simulation environment, and akin to LLMs, we then sample the distribution using the aforementioned probabilities to obtain the next event type. This means every forecast is probabilistic, which means that, for the same context, the output will vary according to the forecasted probabilities.
The features utilized in our LEMs are the following:
  • Event Type: A one-hot encoded categorical feature representing the type of event, such as passes, shots, or fouls, with 33 distinct types considered individually due to their unique distributions related to game context (e.g., location, score).
  • Period: A binary variable indicating the match period (first half = 0, second half = 1), excluding extra time or penalties.
  • Minute: A normalized continuous variable indicating the time of the event occurrence, adjusted to a 0-1 scale by dividing by 60.
  • X and Y: These are continuous features representing the spatial coordinates on the soccer pitch, where the origin (0,0) is located at the bottom left corner from the perspective of the attacking team. The X-coordinate measures the horizontal location from one team’s goal to another, while the Y-coordinate measures the vertical location. Both coordinates are normalized to a range from 0 to 1, ensuring a unified scale regardless of the actual pitch size. This system ensures that actions are consistently recorded from the perspective of the attacking team, thereby standardizing event locations across different matches.
  • IsHomeTeam: A binary indicator denoting whether the event was performed by the home team (1) or the away team (0).
  • IsAccurate: A binary indicator of whether the event was executed accurately (1) or not (0).
  • IsGoal: A binary indicator set to 1 if the event resulted in a goal, otherwise 0.
  • HomeScore and AwayScore: Normalized continuous variables representing the scores of the home and away teams at the event time, scaled to the interval [0,1] by dividing by 10.
The LEM forecasts the same variables as the inputs, with some exceptions: (1) the model forecasts the TimeElapsed variable, which quantifies the temporal distance between consecutive events, rather than predicting Period and Minute, and (2) HomeScore and AwayScore are excluded from prediction, as these scores are directly influenced by the outcomes of other variables that the model forecasts (i.e. if an event leads to a goal, the score changes deterministically).
To generate insights, LEMs are used as a simulation environment. Following the generation of predictions, an interpretation function updates the game state variables (i.e., the context). The new context is then used to generate the next event. This process can be repeatedly applied to simulate the sequence of events throughout a game.
@@
LEMs have advantages and limitations. The positives include their foundational role in creating various metrics for soccer players, ranging from general performance in a season to specific impacts on aspects like shooting quality in particular zones. Simulating numerous games also allows for extracting patterns in the data, enabling comparison with baseline or affected patterns to create insightful metrics [15]. However, a significant limitation is that the model simulates generic football matches, as it was trained on data from all teams, resulting in average game simulations. In Section 3, we will propose a methodology to enable context-specific insights by fine-tuning the models to replicate the behavior of specific teams and players to understand how changing a player’s context affects their performance.
@@ 3.1 Data
This study relies on the Public Wyscout dataset [10] as the primary source of football event data. This dataset encompasses fixtures from top-tier leagues, including the Premier League (England), La Liga (Spanish), Serie A (Italy), Bundesliga (Germany), and Ligue 1 (France), with a focus on the 2017/2018 season. It provides an extensive range of events, from passes and shots to aerial duels and saves. We opted for this dataset since it was the most extensive available at the start of the project. Although the size of the dataset limits the comprehensiveness of the study, it was important for us to provide an open platform to incentive collaboration for further developments.
@@
3.2 Fine Tuning LEMs to Learn Different Contexts
Fine-tuning means adjusting a pre-trained model on a specific task or dataset to improve performance. It involves further training a model already trained on a large dataset, typically on a more specialized or smaller dataset related to the task [12]. The overall process can be seen in Figure 2.
Figure 2: This figure depicts the two-stage process of developing a fine-tuned LEM model: first, we use a large dataset to build a LEM, then we fine-tune it using specific data about our target.
To learn new contexts, we fine-tune the models to reflect the specific behaviors of the teams. In essence, we change the general context-agnostic LEMs to perform another task. Instead of forecasting the next event in an average game of soccer, it will forecast the next event in a game within a predefined context. For this, we select the subset of data that is of interest. For example, if we want to build a model replicating the behavior of one team playing at home, we fine-tune the model using data from that specific team playing at home. Table 2 presents the different types of contexts that we can learn in LEMs. For each type of fine-tuning, we reshape the training data to include exclusively some subsets of data, marked as "Includes". Sometimes overlap occurs. For example in player replacement, when including team data, it will necessarily include data from the player being replaced. Therefore, if we want to remove the effect of the player currently on the team that is going to be replaced by our target, we need to remove all data points related to the outgoing player.
Fine-Tuning Team Opponent Player Player
@@ 3.2 Fine Tuning LEMs to Learn Different Contexts
Table 2: The types of fine-tuning that we can perform on LEMs. The type "Team" focuses on replicating a single team’s behavior against the league average. The type "Player" focuses on estimating the impact of the player on the average team of the league. "Player Addition" includes data from a new player in the context of the new team. "Player Replacement" does the same but excludes data from the player being replaced. In the "Team Face-off" type, we have the data of two teams: the home team and the away team.
This paper focuses on the first four types of fine-tuned LEMs. We have excluded the "Team Face-off" model from this analysis. The rationale behind this exclusion is that "Team Face-off" primarily centers on evaluating potential tactical outcomes of two opposing teams rather than providing insights into player recruitment, which is the primary focus of this study.
3.3 Player Addition vs Player Replacement
Two categories can be used to measure a player’s performance within a new context: "Player Addition" and "Player Replacement." The difference between both approaches is that Player Replacement actively removes the data from a competing player in the squad. On top of adding the influence of the incoming player, we remove the existing influence of the previous player on our squad. Therefore, instead of the training process slowly overwriting the data from the players, it is removed, allowing the training process to start from a blank slate. Table 3 shows the players chosen to be replaced. The criterion for selection of these players was "the attacking player with the least amount of game time from the top 11 most used players in the season".
@@ 3.3 Player Addition vs Player Replacement
Table 3: The list of players that are replaced in the Player Replacement fine-tuning.
3.4 Parameter Tuning
We replicate the procedure to build the general LEM from Mendes-Neves et al. (2024) [9], described in Table 4.
Table 4: Parameters used to train the general LEM.
Model In Out Layers L. Rate Batch Activ.
Type 42 33 [256] 0.0010 32 sigmoid
Accuracy 75 2 [128] 0.0410 1024 sigmoid
Data 77 264 [64, 256, 256] 0.0063 1024 relu
@@ 3.4 Parameter Tuning
$\text{Batch Size}=\log(n^{2}),\quad\text{where}\quad 32\leq\text{Batch Size}\leq 256$ (1)
However, we recommend optimizing these parameters through Bayesian optimization methods (same as the original LEM). We did not incur on this process due to our computational constraints.
The definition of the batch size is the most critical parameter in this work. Fine-tuning a model requires finding the right balance between frequent updates through backpropagation and the training speed of the models [12]. Iteratively, we found Equation 1 to provide the right balance for our use case. The other relevant parameter we must set is the number of simulations for each LEM. As visible in Figure 3, the larger the number of simulations, the lower the variance of the results. We opted for a value of 2500 simulations since it finds a reasonable compromise between simulation time and low variance in the results. Furthermore, we train 10 iterations for each model, allowing us to evaluate the variability of the training process.
Figure 3: This figure illustrates that the outcome variability originates from the model’s training and fine-tuning rather than the simulation process. At the threshold of 3000 simulations, a sequence of 10 consecutive wins is required to alter the expected points by a mere +0.01. Additionally, the figure shows that the error distribution follows a normal curve, indicating that, over an extended period, the average error in simulations is expected to converge towards zero.
3.5 Limitations